Case Inc. is a construction company specializing in custom patios. The patios are constructed of concrete, brick, fiberglass,Journalize the June transactions: (1) for purchase of raw materials, factory labor costs incurred, and manufacturing overheadPost the entries to Work in Process Inventory. (Post the entries to Work in Process Inventory in the order presented in the pPlease help me answer this. Not sure if any of the blanks are right
Answer #1
Journal entries:
Date Acc Titles Debit $ Credit $
1) Raw Mat. Inventory 5100
Accounts Payable 5100
(purchase of RM)
Payroll 5100
Cash 5100
(labor paid off)
Manufacturing overhead 1275
Depreciation exp. 900
Accounts payable 375
(overhead incurred)
2) WIP Inventory 5100
Manuf overhead 1600
R Material 5100
(RM applied to production)
WIP Inventory 3800
Manuf overhead 1300
Payroll 5100
(DL applied to production)
WIP Inventory 4750 (3800*1.25)
Manuf. Overhead 4750
(overhead applied to production)
3) Finished Goods inven. 14730
Manufacturing overhead 14730
(completion of jobs)
Cash 18700
Sales revenue 18700
(sales of jobs)
COGS 14730
Finished Goods inven. 14730
(Cost of sales recorded)
Jobs transferred to FG:
Rodgers Stevens Linton Transf to FG
ob 1275 1825 2350
add:DM 1000 500 1400
add:DL 1300 400 1200
add:OH 1625 500 1500
Total cost 5200 3225 6450 14875
WIP INVENTORY A/C
OB 5450 FG 14875
DM 5100 CB 4225
DL 3800
OH APPLIED 4750                   
19100 Total 19100
WIP CB (UNFINISHED JOB): DM $2200 + DL $900 + OH $(900*1.25) =
2200+900+1125 = 4225
Cost of Goods Manufactured :
Amount $
Op WIP 5450
DM USED 5100
DL USED 3800
OH APPLIED 4750
TOTAL MANUF. COSTS 13650
Total WIP 19100
Less: WIP CB 4225
Cost of goods manufactured 14875